An initially stationary block of mass m is on the floor. A force ofmagnitude 0.500mg is then applied at upward angle θ = 20°.What is the magnitude of the acceleration of the block across thefloor if (a)μs = 0.640 and μk = 0.510 and (b)μs =0.410 and μk = 0.340?

Already found Part A. Don't know what equation to use for part B.If µ is involved in the equation, do you use µs orµk?
Answer #1
Hope this helps. Please rate me.

b) Net force = F -μkmg
= (0.5*cos(20)mg) -0.34mg
=0.13mg

0.13mg/m =0.13g

a=(0.13)(9.8m/s^2) =1.274m/s^2